OpenStudy (anonymous):

Solve 2a - 5d = 30 for d.

OpenStudy (anonymous):

d = (30-2a)/-5

OpenStudy (anonymous):

To solve this, we have to isolate the variable d. The first thing we would do is subtract 2a, which makes the equation look like this: \[-5d=30-2a\] The next step is to divide the -5 from both sides. This is how it's done: \[\frac{ -5d}{ -5 } d=\frac{ 30 }{ -5 }\frac{ -2a }{ -5 }\] \[d=-6-\frac{ 2a }{ -5 }\] Now the equation is solved!

OpenStudy (anonymous):

2a-6